OpenStudy (anonymous):

The longest side of a right triangle is 7 less than 4 times the shortest side. The third side measures 12 inches. Find the length of the shortest side. i got 4x-7, x, and 12 but dont know what to do now??

OpenStudy (anonymous):

please helpppp!!!

OpenStudy (anonymous):

h = 4s - 7 h² = s² + 12² => ( 4s - 7)² = s² + 144 Can you continue?

OpenStudy (anonymous):

nooo so lost..

OpenStudy (anonymous):

Do you know the how to calculate the hypotenuse in the right triangle?

OpenStudy (anonymous):

n o p e

OpenStudy (anonymous):

What do you know about right triangle?

OpenStudy (anonymous):

30, 60, 90

OpenStudy (anonymous):

It's the angles? How about the side?

OpenStudy (anonymous):

with all do respect i just need the answer at this point im on a times test online

OpenStudy (zzr0ck3r):

this is not a place to get answers, we are here to help you get the answer yourself.

OpenStudy (anonymous):

@zzr0ck3r Thanks :)

OpenStudy (anonymous):

@michaelshaw Can you expand ( 4x - 7)² =

OpenStudy (anonymous):

yes 16x^2-56x+49

OpenStudy (anonymous):

-5/3 , 27/5

OpenStudy (anonymous):

What's your y?

OpenStudy (anonymous):

x = 5 the shortest side!